# UDC-VIT: A Real-World Video Dataset for Under-Display Cameras

### v1

@@ 4 Comparison with the Existing UDC Datasets
Many synthetic UDC datasets, including VidUDC33K (Liu et al., 2024), formulate the UDC degradation as follows:
@@
where $I_{t}^{D}$ and $I_{t}^{G}$ denote the UDC-degraded and ground-truth frames, respectively. $\gamma$ is the intensity scaling factor, $k_{t}$ refers to the diffraction kernel (i.e., PSF), $n$ is the noise, and $f$ denotes the clamp function for the pixel value saturation.
Ideally, we would like to compare UDC-VIT with two existing UDC video datasets, PexelsUDC (Chen et al., 2023) and VidUDC33K (Liu et al., 2024). However, since PexelsUDC is not publicly available, we use the P-OLED dataset (Zhou et al., 2021) used to create it. Table 1 gives a summary of the nine UDC datasets. The resolution and frame per second (fps) of UDC-VIT are FHD and 60 fps, respectively, following the Samsung Galaxy Z-Fold 5’s specification.
@@
T/P-OLED (Zhou et al., 2021) Image Synthetic 300 $1024\times 2048\times 3$ CVPR ’21
SYNTH (Feng et al., 2021) Image Synthetic 2,376 $800\times 800\times 3$ CVPR ’21
Pseudo-real (Feng et al., 2023) Image Real 6,747 $512\times 512\times 3$ CVPR ’23
UDC-SIT (Ahn et al., 2024) Image Real 2,340 $1792\times 1280\times 4$ NeurIPS ’23
Tan et al. (2023) Image Synthetic 73,000 - - TCSVT ’23
Wang et al. (2024) Image Synthetic 56,126 - - arXiv ’24
PexelsUDC-T/P (Chen et al., 2023) Video Synthetic $160\times 100$ $1280\times 720\times 3$ 25-50 arXiv ’23
($16,000$)
VidUDC33K (Liu et al., 2024) Video Synthetic $677\times 50$ $1920\times 1080\times 3$ AAAI ’24
($33,850$)
UDC-VIT Video Real $647\times 180$ $1900\times 1060\times 3$ 60
($116,460$)
@@
The camera sensor amplifies the desired signal and unwanted noise in low-light conditions. In the UDC setting, where the sensor is beneath the display panel, the transmittance decreases, leading to amplified noise. The camera sensors with QBC, used in the Samsung Galaxy Z-Fold series (related to UDC-VIT) (Samsung Electronics Co., Ltd., 2021; 2022; 2023; 2024) and ZTE Axon series (related to VidUDC33K) (ZTE Corporation, 2020; 2021; 2022), can influence the noise pattern and pixel intensity (Sony, 2014). Thus, simply adding noise and adjusting the intensity scaling values in Equation 2 may not accurately depict real-world noise and transmittance reduction. For example, in the VidUDC33K dataset, the degraded frame’s noise level is somewhat lower than the ground truth, as shown in Figure 4(b). Similarly, the P-OLED dataset, captured in a controlled setting, exhibits unrealistic noise and excessive transmittance decrease, as depicted in Figure 4(a). In contrast, UDC-VIT in Figure 4(c) accurately shows actual transmittance decrease and digital noise resulting from quantizing digital image signals.
@@
Flares.
Conventional lens flares stem from intense light scattering or reflection within an optical system (Dai et al., 2022; 2023). In contrast, UDC flares arise from light diffraction as it passes through the display panel above the digital camera lens. Thus, it is crucial for each frame in the UDC video dataset to precisely depict the unique flare characteristics, including spatially variant flares, light source variant flares, and temporally variant flares. The P-OLED dataset rarely exhibits flares as it captures images displayed on a monitor in a controlled environment (Figure 6(a) and (d)).
Since UDC distortion increases outward from the camera lens center, spatially variant flares manifest within an image (Yoo et al., 2022). Distorted PSFs must be convolved across different image regions to depict this flare distortion accurately. However, VidUDC33K applies the same PSF convolution across all areas using Equation 2, failing to represent spatially variant flares, as illustrated in Figure 6(b) and (e). Conversely, UDC-VIT effectively captures spatially variant flares (Figure 6(c)).
Various light sources, such as artificial (e.g., LED and halogen) and natural light, can alter the spectra, affecting UDC flares’ shapes. However, VidUDC33K fails to depict light source variant flares. As seen in Figure 6(b) and (e), flare shapes remain similar despite different light sources. Conversely, UDC-VIT exhibits diverse flare shapes, as shown in Figure 6(c) and (f) and Figure 6(b).
A notable characteristic of UDC videos is temporally variant flares caused by the camera’s motion when capturing light sources. The motion results in changes in PSFs (Kwon et al., 2021). However, in the VidUDC33K dataset, attempts to simulate PSF changes through inter-frame homography matrix computations using the method proposed by the previous studies  (Babbar & Bajaj, 2022; Liu et al., 2022a; Ye et al., 2021) yield rare temporally variant flares, as shown in Figure 6(a). Moreover, the shape of typical lens flares in ground-truth frames remains unchanged in degraded frames, indicating the failure of PSF convolution to replicate natural sunlight flares. Conversely, UDC-VIT effectively captures temporally varying flares (Figure 6(b)).
@@
UDC-VIT stands out from other datasets by featuring videos tailored for face recognition (FR). Specifically, we focus on a pair-matching task to identify if two face images belong to the same person. Some datasets, such as T-OLED/P-OLED, SYNTH, and VidUDC33K, only include limited human representations, often too small or from unrecognizable angles (Figure 7(f)). Wang et al. (2024) introduce still image datasets for facial expression recognition (FER), predicting emotions like fear, sadness, anger, and neutrality. However, these datasets are generated using a GAN trained on the P-OLED dataset, which does not adequately simulate actual UDC degradation, notably the lack of flare (Figure 7(e)). Additionally, these datasets are not publicly available. Conversely, UDC-VIT prominently features humans in 64.6% of its videos (approved by the Institutional Review Board (IRB)), featuring various motions (e.g., hand waving, thumbs-up, body-swaying, and walking) by 22 carefully selected subjects from different angles (Figure 7(a)-(d)).
Figure 7: UDC-VIT features human motions, including (a) walking, (b) thumbs-up, (c) hand waving, and (d) body swaying. In contrast, Wang et al. (2024)’s synthetic still image datasets for FER do not show the actual UDC degradations, as shown in (e). Moreover, it is not publicly available. VidUDC33K dataset includes humans but is limited to rear views, as shown in (f).
Less meaningful and strange scenes.
The VidUDC33K dataset often presents unrealistic scenarios. As depicted in Figure 8(a), degraded frames lack UDC flares, displaying flares resembling typical lens flares seen in the ground truth frame. Additionally, flares appear in improbable situations in Figure 8(b), (c), (d), (e), and (f). Moreover, some videos in VidUDC33K may not significantly contribute to research, prompting consideration for their relevance, as shown in Figure 8(g). We specifically analyze the theoretical reasons behind these scenes in Appendix B.
Figure 8: Less meaningful and strange videos in VidUDC33K (Liu et al., 2024). (a) Only lens flare is present, excluding UDC flare. (b) Flare in the seawater and on the cigarette. (c) Flare in the stars of the sky. (d) Flare on the bricks. (e) Flare on the car’s side. (f) Flare on the splashing water droplets. (g) Meaningless abstract image.
@@
To assess the alignment quality of paired videos, we use LoFTR (Sun et al., 2021) as a keypoint matcher, following the convention of the previous studies (Ahn et al., 2024; Feng et al., 2023). We compare the Percentage of Correct Keypoints (PCK), representing the ratio of correctly aligned keypoints to the total number. A keypoint pair is correctly aligned if $d<\alpha\times max(H,W)$, where $d$ is the positional difference between a pair of matched keypoints, $\alpha$ is the threshold, and $H$ and $W$ are the frame or image dimensions. We set $\max(H,W)=1024$ for fair comparison across datasets with varying resolutions.
Table 2 compares alignment accuracy across datasets. The synthetic datasets (e.g., T-OLED/P-OLED, SYNTH, and VidUDC33K) do not require an additional alignment process, leading to PCK values near 100%. In contrast, the Pseudo-real dataset using AlignFormer (Feng et al., 2023), attains a PCK value of 58.75% for $\alpha=0.01$. Unlike Pseudo-real, UDC-VIT maintains PCK values near 100%, demonstrating performance comparable to UDC-SIT, which previously led benchmarks.
Table 2: The comparison of PCK values between the datasets. The UDC-VIT dataset showcases the best alignment quality. It has PCK values close to 100% for all values of $\alpha$.
Dataset Type Need PCK
alignment ($\alpha=0.01$) ($\alpha=0.03$) ($\alpha=0.10$)
T-OLED/P-OLED (Zhou et al., 2021) Image 98.11 98.45 99.08
SYNTH (Feng et al., 2021) Image 99.95 99.96 99.99
Pseudo-real (Feng et al., 2023) Image 58.75 95.08 99.93
UDC-SIT (Ahn et al., 2024) Image 97.26 98.56 99.35
VidUDC33K (Liu et al., 2024) Video 99.82 99.84 99.90
UDC-VIT Video 98.95 99.32 99.69
5 Experiments
This section compares the UDC video restoration performance and face recognition accuracy of existing deep learning models trained by UDC-VIT and the previous synthetic video dataset. Please see Appendix C and D for detailed experiments on shooting conditions and cross-dataset validation.
5.1 Effects on Learnable Restoration Models
In this section, we evaluate the UDC-VIT dataset’s effectiveness by comparing the video restoration performance of six deep learning models on UDC-VIT and VidUDC33K datasets (Liu et al., 2024). The comparison is performed only with VidUDC33K since PexelsUDC is not publicly available. DDRNet (Liu et al., 2024) is the only existing UDC video restoration model, while FastDVDNet (Tassano et al., 2020), EDVR (Wang et al., 2019), and ESTRNN (Zhong et al., 2020) are video restoration models for other general tasks (e.g., deblur, denoising, and super-resolution). DISCNet (Feng et al., 2021) and UDC-UNet (Liu et al., 2022b) are UDC still image restoration models.
Table 3 shows the restoration performance of the six models on both VidUDC33K and UDC-VIT. Interestingly, the performance rankings of the benchmark models across the two datasets do not consistently align. The varying severity of flares between the two datasets is the main reason for the inconsistent restoration performance rankings. Unlike UDC-VIT, VidUDC33K lacks accurate depictions of real-world flares. Examination of input PSNR, SSIM, and LPIPS metrics indicates that their performance degradation on UDC-VIT is more severe than on VidUDC33K. The top performers on UDC-VIT, UDC-UNet and ESTRNN, use residual CNNs to manage complex degradations and enhance restoration quality. They also provide better frame-to-frame consistency than the others, which is crucial for reducing flicker, although some flicker persists. This shows the benefits of residual connections in improving consistency. Note that the restored video of VidUDC33K by DDRNet using their pre-trained model does not create flickering. This result underscores the necessity for research dedicated to UDC’s video restoration using real-world UDC video datasets, an area where UDC-VIT holds promise for significant contributions. Extensive analyses and visual comparisons are available in Appendix C and on our project site.
@@ 5.1 Effects on Learnable Restoration Models
Runtime Param VidUDC33K (Liu et al., 2024) UDC-VIT
@@
DISCNet (Feng et al., 2021) 0.73 3.80 28.89 0.8405 0.2432 24.70 0.8403 0.2675
UDC-UNet (Liu et al., 2022b) 0.37 5.70 28.37 0.8361 0.2561 28.00 0.8911 0.1779
FastDVDNet (Tassano et al., 2020) 0.45 2.48 28.95 0.8638 0.2203 23.89 0.8439 0.2662
EDVR (Wang et al., 2019) 1.17 23.6 28.71 0.8531 0.2416 23.55 0.8331 0.2673
ESTRNN (Zhong et al., 2020) 0.20 2.47 29.54 0.8744 0.2170 25.38 0.8654 0.2216
DDRNet (Liu et al., 2024) 0.44 5.76 31.91 0.9313 0.1306 24.68 0.8539 0.2218
@@ 5.2 Face Recognition
The face recognition (FR) task verifies whether two images are of the same person, similar to typical smartphone applications like Face ID. As shown in Figure 9, we assess average FR accuracy using seven FR models from the DeepFace library (Serengil, 2022), such as VGG-Face (Parkhi et al., 2015), Facenet (Schroff et al., 2015), OpenFace (Baltrušaitis et al., 2016), DeepFace (Taigman et al., 2014), DeepID (Sun et al., 2014), Dlib (King, 2009), and ArcFace (Deng et al., 2019). We test 600 FR frame pairs (human 1 and human 2 from different videos) on a balanced dataset, with 49.2% of the same person (human 1 = human 2) and 50.8% of different people (human 1 $\not$= human 2).
As shown in Figure 9, we compare the effect of human 2’s restoration level in terms of PSNR, SSIM, and LPIPS ($X$-axis) on FR accuracy ($Y$-axis). When evaluating FR accuracy, human 1 is always ground truth (GT), and human 2 can be Input, Restored, or GT. Therefore, Input, Restored, or GT in Figure 9 indicates the group to which human 2 belongs. For example, in Figure 9(a), Input’s PSNR is calculated between human 2 (Input) and human 2 (GT). Similarly, Input’s FR accuracy is calculated between human 1 (GT) and human 2 (Input). To verify the relationship between restoration level and FR accuracy, we illustrate six deep-learning models’ restoration performance (highlighted with a green circle) and corresponding FR accuracy in Figure 9. Restored’s PSNR is calculated between human 2 (Restored) and human 2 (GT). Similarly, Restored’s FR accuracy is calculated between human 1 (GT) and human 2 (Restored).
The results show the significance of leveraging the UDC degradation by deep-learning restoration models to enhance FR accuracy. For example, as depicted in Figure 9(a), Input with PSNR of 16.31 shows 64.5% FR accuracy, UDC-UNet with PSNR of 27.74 shows 82.2% FR accuracy, and GT shows 90.3% FR accuracy.

### v2

@@ 4 Comparison with the Existing UDC Datasets
Many synthetic UDC datasets, including VidUDC33K [31], formulate the UDC degradation as follows:
@@
where $I_{t}^{D}$ and $I_{t}^{G}$ are the UDC-degraded and ground-truth frames, respectively. $\gamma$ is the intensity scaling factor, $k_{t}$ refers to the diffraction kernel (i.e., PSF), $n$ is the noise, and $f$ denotes the clamp function for the pixel value saturation.
Ideally, we would like to compare UDC-VIT with two existing UDC video datasets, PexelsUDC-T/P [12] and VidUDC33K [31]. However, since PexelsUDC-T/P is not publicly available, we use the P-OLED dataset [63], which is used to train the GAN model that generates synthetic PexelsUDC-T/P videos. Tab. 1 gives a summary of the nine UDC datasets. The resolution and frame rate (fps) of UDC-VIT are FHD and 60 fps, respectively, in line with the Samsung Galaxy Z-Fold 5’s specifications.
@@
T/P-OLED [63] Image Synthetic 300 1024×2048×31024$\times$ 2048$\times$ 31024 × 2048 × 3 - CVPR ’21
SYNTH [18] Image Synthetic 2,376 800×800×3800$\times$ 800$\times$ 3800 × 800 × 3 - CVPR ’21
Pseudo-real [20] Image Real 6,747 512×512×3512$\times$ 512$\times$ 3512 × 512 × 3 - CVPR ’23
UDC-SIT [1] Image Real 2,340 1792×1280×41792$\times$ 1280$\times$ 41792 × 1280 × 4 - NeurIPS ’23
Tan et al[53] Image Synthetic 73,000 - - TCSVT ’23
Wang et al[58] Image Synthetic 56,126 - - ICIP ’24
PexelsUDC-T/P [12] Video Synthetic 160×100160$\times$ 100160 × 100 1280×720×31280$\times$ 720$\times$ 31280 × 720 × 3 25-50 arXiv ’23
(16,00016,00016 , 000)
VidUDC33K [31] Video Synthetic 677×50677$\times$ 50677 × 50 1920×1080×31920$\times$ 1080$\times$ 31920 × 1080 × 3 - AAAI ’24
(33,85033,85033 , 850)
UDC-VIT Video Real 647×180647$\times$ 180647 × 180 1900×1060×31900$\times$ 1060$\times$ 31900 × 1060 × 3 60 ICCV ’25
(116,460116,460116 , 460)
@@
The camera sensor amplifies the desired signal and unwanted noise in low-light conditions. In the UDC setting, where the sensor is beneath the display panel, the transmittance decreases, leading to amplified noise. The camera sensors with QBC, used in the Samsung Galaxy Z-Fold series (related to UDC-VIT) [43, 44, 45, 46] and ZTE Axon series (related to VidUDC33K) [64, 65, 66], can influence the noise pattern and pixel intensity [49]. Thus, simply adding noise and adjusting the intensity scaling values in Eq. 2 may not accurately depict real-world noise and transmittance reduction. For example, in the VidUDC33K dataset, the degraded frame’s noise level is somewhat lower than the ground truth, as shown in Fig. 4(b). Similarly, the P-OLED dataset, captured in a controlled setting, exhibits unrealistic noise and excessive transmittance decrease, as depicted in Fig. 4(a). In contrast, UDC-VIT in Fig. 4(c) accurately shows actual transmittance decrease and digital noise resulting from quantizing digital image signals.
@@
Flares.
Conventional lens flares stem from intense light scattering or reflection within an optical system [14, 15]. In contrast, UDC flares arise from light diffraction as it passes through the display panel above the digital camera lens. Thus, it is crucial for each frame in the UDC video dataset to precisely depict the unique flare characteristics, including spatially variant flares, light source variant flares, and temporally variant flares. The P-OLED dataset rarely exhibits flares as it captures images displayed on a monitor in a controlled environment (Fig. 5(a) and (d)).
Since UDC distortion increases outward from the camera lens center, spatially variant flares manifest within an image [61]. Distorted PSFs must be convolved across different image regions to depict this flare distortion accurately. However, VidUDC33K applies the same PSF convolution across all areas using Eq. 2, failing to represent spatially variant flares, as illustrated in Fig. 5(b) and (e). Conversely, UDC-VIT effectively captures spatially variant flares (Fig. 5(c)).
Various light sources, such as artificial (e.g., LED and halogen) and natural light, can alter the spectra, affecting UDC flares’ shapes. However, VidUDC33K fails to depict light source variant flares. As seen in Fig. 5(b) and (e), flare shapes remain similar despite different light sources. Conversely, UDC-VIT exhibits diverse flare shapes, as shown in Fig. 5(c) and (f) and Fig. 6(b).
A notable characteristic of UDC videos is temporally variant flares caused by the camera’s motion when capturing light sources. The motion results in changes in PSFs [29]. However, in the VidUDC33K dataset, attempts to simulate PSF changes through inter-frame homography matrix computations using the method proposed by the previous studies [9, 32, 60] yield rare temporally variant flares, as shown in Fig. 6(a). Moreover, the shape of typical lens flares in ground-truth frames remains unchanged in degraded frames, indicating the failure of PSF convolution to replicate natural sunlight flares. Conversely, UDC-VIT effectively captures temporally varying flares (Fig. 6(b)).
@@
UDC-VIT stands out from other datasets by featuring videos tailored for face recognition (FR). Specifically, we focus on a pair-matching task to identify if two face images belong to the same person. Some datasets, such as T-OLED/P-OLED, SYNTH, and VidUDC33K, only include limited human representations, often too small or from unrecognizable angles (Fig. 7(f)). Wang et al[58] introduce still image datasets for facial expression recognition (FER), predicting emotions like fear, sadness, anger, and neutrality. However, these datasets are generated using a GAN trained on the P-OLED dataset, which does not adequately simulate actual UDC degradation, notably the lack of flare (Fig. 7(e)). Additionally, these datasets are not publicly available. Conversely, UDC-VIT prominently features humans in 64.6% of its videos (approved by the Institutional Review Board (IRB)), featuring various motions (e.g., hand waving, thumbs-up, body-swaying, and walking) by 22 carefully selected subjects from different angles (Fig. 7(a)-(d)).
Figure 7: UDC-VIT features human motions, including (a) walking, (b) thumbs-up, (c) hand waving, and (d) body swaying. In contrast, Wang et al[58]’s synthetic still image datasets for FER do not show the actual UDC degradations, as shown in (e). Moreover, it is not publicly available. VidUDC33K dataset includes humans but is limited to rear views, as shown in (f).
Figure 8: Strange scenes and white artifacts in VidUDC33K [31]. (a) Flare in the seawater and on the cigarette. (b) Flare on the bird’s beak. (c) Flare on fruit. (d) Only lens flare is present, excluding UDC flare. (e) Black frame resulting from an incorrectly transformed PSF. (f) Washed-out colors on clouds. (g) Washed-out colors on buildings.
Strange scenes and white artifacts.
The VidUDC33K dataset includes degraded frames with implausible flare artifacts, often appearing in unrelated regions, such as seawater, bird’s beak, or fruits, as depicted in Fig. 8(a)-(c). These artifacts arise from flawed synthesis using PSF convolution after pixel scaling, which indiscriminately amplifies bright areas regardless of actual light sources. Some videos also lack UDC-specific flares altogether, instead presenting only natural lens flares or semantically empty scenes (Fig. 8(d)-(e)). In addition, excessive scaling introduces white artifacts that wash out natural colors in bright areas, such as clouds or white walls (Fig. 8(f)-(g)), raising concerns about the realism and practical utility of such data. A theoretical analysis appears in Appendix A.
[…]
Dataset Align Align PCK0.003 PCK0.01 PCK0.03 PCK0.10
Req. method
T/P-OLED [63] 96.32 98.11 98.45 99.08
SYNTH [18] 99.70 99.95 99.96 99.99
Pseudo-real [20] AlignFormer [20] N/A 58.75 95.08 99.93
UDC-SIT [1] DFT 93.67 97.26 98.56 99.35
VidUDC33K [31] 99.65 99.82 99.84 99.90
UDC-VIT 85.10 98.65 99.22 99.64
UDC-VIT DFT 92.12 98.95 99.32 99.69
@@
To assess the alignment quality of paired videos, we use LoFTR [50] as a keypoint matcher, following the convention of the previous studies [1, 20]. We compare the Percentage of Correct Keypoints (PCK), representing the ratio of correctly aligned keypoints to the total number. A keypoint pair is correctly aligned if $d<\alpha\times max(H,W)$, where $d$ is the positional difference between a pair of matched keypoints, $\alpha$ is the threshold, and $H$ and $W$ are the frame or image dimensions. We set $max(H,W)=1024$ for fair comparison across datasets with varying resolutions.
Tab. 2 compares alignment accuracy across datasets. The synthetic datasets (e.g., T-OLED/P-OLED, SYNTH, and VidUDC33K) do not require an additional alignment process, leading to PCK values near 100%. In contrast, the Pseudo-real dataset using AlignFormer [20], attains a PCK value of 58.75% for $\alpha=0.01$. Unlike Pseudo-real, UDC-VIT achieves consistently high PCK values (92.12–99.69%), demonstrating performance comparable to UDC-SIT, which previously led benchmarks.
@@
Runtime Param VidUDC33K [31] UDC-VIT
@@
DISCNet [18] 0.73 3.80 28.89 0.8405 0.2432 24.70 0.8403 0.2675
UDC-UNet [33] 0.37 5.70 28.37 0.8361 0.2561 28.00 0.8911 0.1779
FastDVDNet [54] 0.45 2.48 28.95 0.8638 0.2203 23.89 0.8439 0.2662
EDVR [57] 1.17 23.6 28.71 0.8531 0.2416 23.55 0.8331 0.2673
ESTRNN [62] 0.20 2.47 29.54 0.8744 0.2170 25.38 0.8654 0.2216
DDRNet [31] 0.44 5.76 31.91 0.9313 0.1306 24.68 0.8539 0.2218
5 Experiments
This section compares the UDC video restoration performance and face recognition accuracy of existing deep learning models trained by UDC-VIT and the previous synthetic video dataset. Please see Appendix B and C, as well as https://mcrl.github.io/UDC for extensive analyses and visual comparisons, including cross-dataset validation, shooting conditions, and flickering artifacts in restored videos.
5.1 Effects on Learnable Restoration Models
In this section, we evaluate the effectiveness of the UDC-VIT dataset by comparing the video restoration performance of six deep learning models, each trained and tested on the UDC-VIT and VidUDC33K datasets [31]. The comparison is performed only with VidUDC33K since PexelsUDC-T/P is not publicly available. DDRNet [31] is the only existing UDC video restoration model, while FastDVDNet [54], EDVR [57], and ESTRNN [62] are video restoration models for other general tasks (e.g., deblur, denoising, and super-resolution). DISCNet [18] and UDC-UNet [33] are UDC still image restoration models.
Tab. 3 shows the restoration performance of the six models on both VidUDC33K and UDC-VIT. Interestingly, the performance rankings of the benchmark models across the two datasets do not consistently align. The varying severity of flares between the two datasets is the main reason for the inconsistent restoration performance rankings. Unlike UDC-VIT, VidUDC33K does not accurately represent real-world degradations (e.g., flares), as reflected in its higher input-level PSNR and SSIM, and lower LPIPS. The top-performing models on UDC-VIT, such as UDC-UNet and ESTRNN, use residual CNNs to address complex degradations and improve restoration quality. They also provide better frame-to-frame consistency than the others, which is crucial for reducing flicker, although some flicker persists. This shows the benefits of residual connections in improving consistency. Note that the restored video of VidUDC33K by DDRNet, using their pre-trained model, does not exhibit flickering. This highlights the need for real-world UDC video datasets, such as UDC-VIT, and is also evident in cross-dataset experiments presented in Appendix B, which demonstrate the limitations of synthetic data in modeling complex degradations.
@@ 5.2 Face Recognition
The face recognition (FR) task verifies whether two images are of the same person, similar to typical smartphone applications like Face ID. As shown in Fig. 9, we assess average FR accuracy using seven FR models from the DeepFace library [48], such as VGG-Face [37], Facenet [47], OpenFace [10], DeepFace [52], DeepID [51], Dlib [28], and ArcFace [16]. We test 600 FR frame pairs (human 1 and human 2 from different videos) on a balanced dataset, with 49.2% of the same person (human 1 = human 2) and 50.8% of different people (human 1 $\not$= human 2). Note that Tab. 3 ranks models on the complete UDC-VIT test set, while Fig. 9 evaluates only a sampled subset of the face-related data.
As shown in Fig. 9, we compare the effect of human 2’s restoration level in terms of PSNR, SSIM, and LPIPS ($X$-axis) on FR accuracy ($Y$-axis). When evaluating FR accuracy, human 1 is always ground truth (GT), and human 2 can be Input, Restored, or GT. Therefore, Input, Restored, or GT in Fig. 9 indicates the group to which human 2 belongs. For example, in Fig. 9(a), Input’s PSNR is calculated between human 2 (Input) and human 2 (GT). Similarly, Input’s FR accuracy is calculated between human 1 (GT) and human 2 (Input). To verify the relationship between restoration level and FR accuracy, we illustrate six deep-learning models’ restoration performance (highlighted with a green circle) and corresponding FR accuracy in Fig. 9. Restored’s PSNR is calculated between human 2 (Restored) and human 2 (GT). Similarly, Restored’s FR accuracy is calculated between human 1 (GT) and human 2 (Restored).
@@
The results demonstrate the importance of leveraging UDC degradation by deep-learning restoration models to improve FR accuracy, consistent with prior work [3]. For example, as depicted in Fig. 9(a), Input with PSNR of 16.31 shows 64.5% FR accuracy, UDC-UNet with PSNR of 27.74 shows 82.2% FR accuracy, and GT shows 90.3% FR accuracy.